OpenStudy (anonymous):

integral of 3x^2 4sprt of 9+3x^3 dx

OpenStudy (amistre64):

depending on notation .... try a by parts run; poly go to zero rather easily

hartnn (hartnn):

do a substitution of u=9+3x^3 then du =.... ? (in your question,combine the x^2 with dx)

OpenStudy (loser66):

\[\huge \int {3x^2~\sqrt[4]{9+x^3}dx}\] is it ?

OpenStudy (amistre64):

that does make it easier to assess :)

OpenStudy (anonymous):

you could write the integral as integral of (9+x^3)^1/4 dx^3

OpenStudy (anonymous):

Try u substitution. \[u = 9+ x ^{3}\] \[\frac{ du }{ dx } = 3x ^{2}\] \[du = 3x^{2} dx\] then \[\int\limits_{?}^{?} \sqrt[4]{u} du\]